# Discrete Curvature on Graphs

### v1

@@ 3.2 Construction of the inner parallel bodies
The next step of the proof constructs the inner parallel bodies of both $K$ and its Thick Convex Hull.
Definition 7 (Inner parallel bodies $K_{-\rho}$ of $K$).
@@
where $B_{\rho}(x)$ is a ball of radius $\rho$ centered at $x$. The greatest number $\rho$ for which $K_{-\rho}$ is not empty is the inradius $in(K)$ of $K$. Equivalently, $in(K)$ is the radius of the largest ball contained in $K$.
Remark: Note that for all $N\in\mathbb{N}$, there exists a strictly convex body body $\Omega\subset\mathbb{H}^{2}$ such that it has $N$ inscribed balls not connected by continuous motion within $\Omega$. We can construct this body $H$ by taking of the convex hull of $N$ equal balls, suitably choosen, and considering the outer parallel body $H_{\epsilon}$ for a small $\epsilon>0$. This ensures that $H_{\epsilon}$ is strictly convex. This phenomena is not encountered in the Euclidean space, where there exist a unique inscribed ball for a strictly convex body $\Omega\subset\mathbb{R}^{2}$.
Remark: Note that $K$ being convex does not guarantee that $K_{-\rho}$ is convex, as shown in Figure 10. In that case, $K$ is the convex hull of two balls $B^{1}$ and $B^{2}$ of radius $R$. The inner parallel body $K_{-\rho}$ for $0<\rho<in(K)$ is the concatenation of the hypercircles a distance $\rho$ from the geodesics with the two balls of radius $r=R-\rho$. It is well known that hypercircles are convex, therefore, there exist a geodesic $[x,y]$ for $x\in B^{1}_{-\rho}$ and $y\in B^{2}_{-\rho}$ such that $[x,y]\nsubseteq K_{-\rho}$ (represented by the orange line in Figure 10 below). This fact is a consequence of the previous remark.
Figure 10: Example of an inner body of a convex body which is not convex
Because of this Remark, is not immediate that the inner parallel body of the Thick Convex Hull of $K$ is convex. We will proceed to prove this is indeed the case with the following lemmas.
@@
Let $S$ and $K$ be convex subsets of $\mathbb{H}^{2}$. If $S\subset K$, then $S_{-\rho}\subset K_{-\rho}$ for $0\leq\rho\leq in(S)$.
@@
Let $x\in S_{-\rho}$. By the definition of inner parallel body $B_{\rho}(x)\subset S\subset K$, which implies that $x\in K_{-\rho}$. This is satisfied $\forall x\in S_{-\rho}$, so we conclude that $S_{-\rho}\subset K_{-\rho}$, as required. ∎
Combining Theorem 1 and Lemma 3.5, we get the following inclusion relationships:
ThickConv ($\text{ThickConv }(T,\lambda)\subset K$
$\left(\text{ThickConv }(T,\lambda)\right)_{-\rho}\subset K_{-\rho}$
Theorem 2.
Let $K$ be a thick $\lambda$-concave body, let $T$ be the cut locus of $K$ and $H=\text{ThickConv }(T,\lambda)$. Then $H_{-\rho}$ is a convex simply connected domain for $0\leq\rho\leq\coth^{-1}(\lambda)$.
@@
We will first prove that $H_{-\rho}$ is simply connected. Let us define an arbitrary ordering of the vertices of $T$ such that
ThickConv ($\text{ThickConv }(T,\lambda)=W\cup[B_{\lambda}(x_{i}),B_{\lambda}(x_{i+1})]_{\lambda},$
where $W$ is the convex hull of $\bigcup B_{i}$, as of definition 6, and $[B_{\lambda}(x_{i}),B_{\lambda}(x_{i+1})]_{\lambda}$ is the region bounded by arcs of hypercircle of curvature $1/\lambda$ and two circular arcs of curvature $\lambda$ centered at two consecutive $x_{i}$ and $x_{i+1}$, as shown in Figure 11. Then, by construction $H_{-\rho}$ is simply connected and well-defined for $0\leq\rho\leq 1/c\cdot\coth^{-1}(\lambda/c)$.
Figure 11: Construction of $H$ and $H_{-\rho}$ from the proof of Theorem 2 for vertices $x_{1}$ and $x_{2}$
Now, if $H_{-\rho}$ is a simply connected domain and $\partial H_{-\rho}$ is $C^{1}-$smooth with $\kappa(x)\geq 0\,\,\forall x\in\partial H_{-\rho}$ (consistent curvature sign in the boundary), then $H_{-\rho}$ is convex (proof of this statement can be found in [21]). ∎
@@ 3.3 Steiner’s formula in the Hyperbolic space
We will now use Steiner’s formula to relate the area of a body $K$ with the area of its inner parallel body. According to Steiner’s classical result, if $\partial K$ is an arbitrary convex curve in the plane of length $|\partial K|$ and area $|K|$ and $\partial K_{\rho}$ is a curve parallel to $\partial K$ at a distance $\rho$ with length $|\partial K_{\rho}|$ and area $|K_{\rho}|$. Then,
$|\partial K_{\rho}|=|\partial K|+2\pi\rho,\quad|K_{\rho}|=|K|+\rho|\partial K|+\pi\rho^{2}$
We would like to work with a generalization of the Steiner’s formula for the Hyperbolic space. To do that, we have to choose a coordinate system in which $v^{1}=0$ is the curve $\partial K$, and in which $v^{2}=\text{constant}$ are the geodesics orthogonal to $\partial K$, where $\partial K$ is a simple, closed, bounding, and differentiable curve. Further let $v^{2}$ be the arc length of $\partial K$ measured positively for motion on the curve which keeps the bounded area to the left, and let $v^{1}$ be the arc length of geodesics normal to $\partial K$ measured positively outward from $\partial K$. Choose the unit normals to $\partial K$ so that they point toward the interior of $\partial K$. We can now use the following Lemma due to Steiner’s [22]:
Lemma 3.6.
(Steiner’s formula in spaces of negative curvature [22]) Choose $\partial K$ and the coordinate system as above, satisfying
$\kappa_{g}(v^{2})\sinh(v^{1})+\cosh(v^{1})>0$ (1)
for $\kappa_{g}(v^{2})>0$ and $\rho\geq 0$. The area of the inner parallel body $A_{\rho}$ with respect to $A$ is then
$|K_{\rho}|=|\partial K|\sinh\left(\rho\right)-2\pi\left(\cosh\left(\rho\right)-1\right)+|K|\cosh\left(\rho\right).$ (2)
For our purpose, we want to work with the Steiner inequality when $\rho\leq 0$. Setting $v^{1}=-\rho$ for $\rho\geq 0$, we have that inequality 1 can then be adapted as follows:
$\kappa_{g}(v^{2})<\coth(\rho)$ (3)
and formula 2 becomes
$|K_{-\rho}|=-|\partial K|\sinh\left(\rho\right)-2\pi\left(\cosh\left(\rho\right)-1\right)+|K|\cosh\left(\rho\right)$
Rearranging, we get that
$|K|=|K_{-\rho}|\operatorname{sech}(\rho)+|\partial K|\tanh\left(\rho\right)+2\pi\left(1-\operatorname{sech}(\rho)\right).$ (4)
Note that inequality 3 is included in the assumptions of Theorem A when setting $\kappa<\lambda$.
We will are now ready to finalize the proof of Theorems A and B.
Proof of Theorem A:Let $K\subset\mathbb{H}^{2}$ be a thick $\lambda$-concave body and $S\subset\mathbb{H}^{2}$ be a thick $\lambda$-sausage. For $\rho=\coth^{-1}(\lambda)$, $S_{-\rho}$ is a geodesic with $|S_{-\rho}|=0$. Now, let $T^{\prime}$ be the cut locus of $S$ with $H^{\prime}=\text{ThickConv }(T^{\prime},\lambda)$, and let $T$ be the cut locus of $K$ with $H=\text{ThickConv }(T,\lambda)$. Then, by Theorem 2, $H_{-\rho}\subset K_{-\rho}$ is a simply connected domain, and $0=|S_{-\rho}|=|H^{\prime}_{-\rho}|\leq|H_{-\rho}|\leq|K_{-\rho}|$. Now, if $|\partial S|=|\partial K|$ by Steiner’s result we get that $|S|\leq|K|$ as required.
@@
To proof the other direction, notice that if $|K_{-\rho}|=0$ then $|K_{-\rho}|$ is a convex region with null area and, hence, a geodesic. $K_{-\rho}$ is a a geodesic when $K$ is thick $\lambda$-sausage body, as defined in Definition 6, and $\rho=\coth^{-1}(\lambda)$.
@@
Proof of Theorem B: Substituting $\rho=\coth^{-1}(\lambda)$ in equation (2), we get that the area of the thick $\lambda$-sausage is
$|K|=\frac{|\partial K|}{\lambda}+2\pi\left(1-\sqrt{1-\frac{1}{\lambda^{2}}}\right).$ (5)
Applying the appropriate scaling, when $K\subset\mathbb{H}^{2}(-c^{2})$, (5) becomes
@@
It is easy to check (applying the Hôpital Rule twice) that as $c\rightarrow 0$ from the left, equation 5 approaches the result in the Euclidean space proved in [4].
This concludes the proof of one direction of Theorem B. The other direction is trivial.
Proof of Theorem C: We will prove Theorem C by contradiction. Let $K$ be a thick $\lambda$-concave body and assume a ball $B$ of curvature $\lambda$ can not roll freely in $K$. Then, there exists a ball $B^{\prime}$ of curvature $\lambda^{\prime}>\lambda$ such that $B^{\prime}$ touches $\partial K$ at least twice. Name two of these points of intersection $x$ and $y$. Now, let $z$ be the center of $B^{\prime}$ and consider the geodesic $[x,z]$ and $[y,z]$, as shown in Figure 12. By construction, these geodesics are orthogonal to $\partial K$ so, by Definition 3, $z$ belongs to the cut locus $T$ of $K$. Hence, the distance $\text{d}(T,\partial K)\leq\coth^{-1}(\lambda^{\prime})<\coth^{-1}(\lambda)$. Nevertheless, in Lemma 3.5 we proved that ThickConv ($\text{ThickConv }(T,\lambda)\subset K$ and, hence, d($T,\partial K$) $\geq\coth^{-1}(\lambda)$, arriving at a contradiction.
Figure 12: Proof of Theorem C
To prove that the result is false if the curvature is bounded as $1/\lambda^{\prime}\leq k\leq\lambda$ for some $\lambda^{\prime}>\lambda$, consider the convex body $Q$ bounded by two equal circular arcs of curvature $\lambda$ and two equal arcs of hypercircle of curvature $1/\lambda-\epsilon$ for $\epsilon>0$. Then, by construction, a ball of curvature $\lambda$ will not be able to roll freely inside $Q$.

### v2

@@ 3.1 Construction of the inner parallel bodies
The next step of the proof constructs the convex inner parallel bodies of $K$.
Definition 3.1 (Inner parallel bodies $K_{-\rho}$ of $K$).
@@
where $B_{\rho}(x)$ is a ball of radius $\rho$ centered at $x$. The greatest number $\rho$ for which $K_{-\rho}$ is not empty is the inradius, denoted as $in(K)$, of $K$. Equivalently, $in(K)$ is the radius of the largest ball contained in $K$.
@@
Let $L$ and $K$ be convex subsets of $\mathbb{H}^{2}$. If $L\subset K$, then $L_{-\rho}\subset K_{-\rho}$ for $0\leq\rho\leq in(L)$.
@@
Let $x\in L_{-\rho}$. By the definition of inner parallel body $B_{\rho}(x)\subset L\subset K$, which implies that $x\in K_{-\rho}$. This is satisfied $\forall x\in L_{-\rho}$, so we conclude that $L_{-\rho}\subset K_{-\rho}$, as required. ∎
Remark 3.2.
Note that for all $N\in\mathbb{N}$, there exists a strictly convex body $\Omega\subset\mathbb{H}^{2}$ such that it has $N$ inscribed balls not connected by continuous motion within $\Omega$. We can construct this body $H$ by taking of the convex hull of $N$ equal balls, suitably choosen, and considering the outer parallel body $H_{\rho}$ for a small $\rho>0$. This ensures that $H_{\rho}$ is strictly convex. This phenomenon is not encountered in the Euclidean space, where there exist a unique inscribed ball for a strictly convex body $\Omega\subset\mathbb{R}^{2}$.
Remark 3.3.
Note that $K$ being convex does not guarantee that $K_{-\rho}$ is convex, as shown in Figure 6. In that example, $K$ is the convex hull of two balls $B^{1}$ and $B^{2}$ of radius $R$. The inner parallel body $K_{-\rho}$ for $0<\rho<in(K)$ is the domain bounded by the hypercircles a distance $\rho$ from the geodesics and two balls of radius $r=R-\rho$. It is well known that the domain bounded by hypercircles is convex, therefore, there exist a geodesic $[x,y]$ for $x\in B^{1}_{-\rho}$ and $y\in B^{2}_{-\rho}$ such that $x,y]\nsubseteq K_{-\rho}$ (represented by the orange line in Figure 6 below). This phenomenon is a consequence of the previous Remark.
Figure 6: Example of an inner parallel body of a convex body which is not convex
Because of this Remark, it is not immediate that the inner parallel body of $K$ is convex. We will proceed to prove this construction is indeed possible with the following lemma.
Lemma 3.4.
Let $K$ be a thick $\lambda$-concave body. Then, given $0<\rho<in(K)$, there exists a non-empty convex set $L$ such that $K$ is the closed $\rho$-neighborhood of $L$. That is, $K$ is the union of all closed balls $B_{\rho}$ of radius $\rho$ whose centers lie in $L$. We can denote this as the Minkowski sum $K=L+B_{\rho}$.
@@
This is a direct consequence of Theorem 1 in [11], when taking $\epsilon=\tanh^{-1}(\lambda)$. Note that we can apply this result because all the bodies that we are working with have $C^{1,1}$ smooth boundary. ∎
Remark 3.5.
Lemma 3.4 provides a proof of Blaschke’s rolling theorem for thick $\lambda$-concave bodies (see Figure 7). In $\mathbb{R}^{2}$, the statement is as follows: let $\gamma$ be a closed, simple, smooth, convex curve in $\mathbb{R}^{2}$ such that for all $x\in\gamma$, $\kappa_{g}(x)\leq\lambda$. Then a ball of curvature $\lambda$ rolls along the boundary freely inside the region bounded by $\gamma$. This result was extended to $\mathbb{R}^{n}$ by [12, 13], and later to Riemannian manifolds of bounded curvature in [14, 15, 16, 17]. Nevertheless, note that this statement does not hold in general for $\lambda$-concave domains in $\mathbb{H}^{2}$: consider the convex body $Q$ bounded by two equal circular arcs of curvature $\lambda$ and two equal arcs of a hypercircle of curvature $1/\lambda-\epsilon$, with $\epsilon>0$. By construction, a ball of curvature $\lambda$ cannot roll freely inside $Q$ (see Figure 6).
Figure 7: Blaschke’s rolling theorem for thick $\lambda$-concave bodies in $\mathbb{H}^{2}$
@@ 3.2 Steiner’s formula in the Hyperbolic Space
We will now use Steiner’s formula to relate the area of a body $K$ with the area of its inner parallel body. According to Steiner’s classical result, let $\partial K$ be an arbitrary convex curve in the plane of length $|\partial K|$, bounding the area $|K|$. Let $\partial K_{\rho}$ be a curve parallel to $\partial K$ at a distance $\rho$, with length $|\partial K_{\rho}|$ and bounding the area $|K_{\rho}|$. Then,
$|\partial K_{\rho}|=|\partial K|+2\pi\rho,\quad|K_{\rho}|=|K|+\rho|\partial K|+\pi\rho^{2}.$
In [18], the author generalizes these formulae to simple, closed, and differentiable curves lying in spaces of constant curvature and satisfying certain curvature boundary constraints. Nevertheless, we are unable to apply this formula in our context, since $\partial K$ is not necessarily $C^{2}$. In [19, 20], the authors extend these formulae to compact convex sets in the sphere and in Euclidean space using valuations, and later [21] applies the same technique to generalize the result to hyperbolic space.
Lemma 3.6 (Steiner formula for compact convex bodies in the hyperbolic space [21]).
Let $K$ be a compact convex body in $\mathbb{H}^{2}$, and let $\rho\geq 0$, then
$|K_{\rho}|=|\partial K|\sinh\left(\rho\right)+2\pi\left(\cosh\left(\rho\right)-1\right)+|K|\cosh\left(\rho\right).$ (3.1)
Proof.
The proof of this Lemma can be found in Corollary 4.5 of [21]. ∎
Corollary 3.6.1.
Let $K$ be a thick $\lambda$-concave body in $\mathbb{H}^{2}$, and let $0\leq\rho<in(K)$, then
$|K_{-\rho}|=-|\partial K|\sinh\left(\rho\right)+2\pi\left(\cosh\left(\rho\right)-1\right)+|K|\cosh\left(\rho\right).$ (3.2)
and
$|\partial K_{-\rho}|=-\sinh(\rho)\big(|K|+2\pi\big)+|\partial K|\cosh(\rho)$ (3.3)
Proof.
Let $0\leq\rho<in(K)$ and $t>0$. By Lemma 3.6 we have that
$|K_{\rho+t}|$ $=|\partial K|\sinh(\rho+t)+2\pi\big(\cosh(\rho+t)-1\big)+|K|\cosh(\rho+t)$
$=|\partial K_{\rho}|\sinh(t)+2\pi\big(\cosh(t)-1\big)+|K_{\rho}|\cosh(t).$
Rearranging, we get
$|\partial K_{\rho}|\sinh(t)+(|K_{\rho}|+2\pi)\cosh(t)$ $=|\partial K|\big(\cosh(\rho)\,\sinh(t)+\sinh(\rho)\,\cosh(t)\big)$
$\quad+(|K|+2\pi)\big(\cosh(\rho)\,\cosh(t)+\sinh(\rho)\,\sinh(t)\big)$
$=\cosh(t)\Big(\cosh(\rho)(|K|+2\pi)+\sinh(\rho)|\partial K|\Big)$
$\quad+\sinh(t)\Big(\sinh(\rho)(|K|+2\pi)+\cosh(\rho)|\partial K|\Big).$
Since this must hold for all $t$, the terms in front $\cosh(t)$ and $\sinh(t)$ of the formula above must be equal. That is:
$\begin{pmatrix}|K_{\rho}|+2\pi\\[6.0pt] |\partial K_{\rho}|\end{pmatrix}=\begin{pmatrix}\cosh(\rho)\quad&\quad\sinh(\rho)\\[6.0pt] \sinh(\rho)\quad&\quad\cosh(\rho)\end{pmatrix}\begin{pmatrix}|K|+2\pi\\[6.0pt] |\partial K|\end{pmatrix}$.
Rearranging, we get to
$\begin{pmatrix}|K|+2\pi\\[6.0pt] |\partial K|\end{pmatrix}=\begin{pmatrix}\cosh(\rho)&-\sinh(\rho)\\[6.0pt] -\sinh(\rho)&\cosh(\rho)\end{pmatrix}\begin{pmatrix}|K_{\rho}|+2\pi\\[6.0pt] |\partial K_{\rho}|\end{pmatrix}$. (3.4)
Now, from Lemma 3.4 we know that a thick $\lambda$-concave body $K$ is the outer parallel body of some convex body $K_{-\rho}$, for $0\leq\rho<in(K)$ . Hence, expression (3.4) becomes
$\begin{pmatrix}|K_{-\rho}|+2\pi\\[6.0pt] |\partial K_{-\rho}|\end{pmatrix}=\begin{pmatrix}\cosh(\rho)&-\sinh(\rho)\\[6.0pt] -\sinh(\rho)&\cosh(\rho)\end{pmatrix}\begin{pmatrix}|K|+2\pi\\[6.0pt] |\partial K|\end{pmatrix}$ (3.5)
as required. ∎
Note that the result of Corollary 3.6.1 is analogous to the result obtained in [18] for bodies in $\mathbb{H}^{2}$ with $C^{2}$ smooth boundary.
We are now ready to finalize the proof of Theorems 1.2 and 1.3.
Proof of Theorem 1.2: Let $K\subset\mathbb{H}^{2}$ be a thick $\lambda$-concave body and $S\subset\mathbb{H}^{2}$ be a thick $\lambda$-sausage. Rearranging formula (3.2) we get
$|K|=|K_{-\rho}|\operatorname{sech}(\rho)+|\partial K|\tanh\left(\rho\right)+2\pi\left(\operatorname{sech}(\rho)-1\right).$ (3.6)
Let $\rho_{\lambda}=\coth^{-1}(\lambda)$. Then, $S_{-\rho_{\lambda}}$ is a geodesic segment with $|S_{-\rho_{\lambda}}|=0$. Then, $0=|S_{-\rho_{\lambda}}|\leq|K_{-\rho_{\lambda}}|$. By the assumption of Theorem 1.2, we have $|\partial S|=|\partial K|$. Then,by applying formula (3.6), it follows that $|S|\leq|K|$, as required.
@@
To prove the other direction, notice that if $|K_{-\rho_{\lambda}}|=0$ then $|K_{-\rho_{\lambda}}|$ is a convex region with null area and, hence, a geodesic segment. Moreover, $K_{-\rho_{\lambda}}$ is a geodesic segment whenever $K$ is a thick $\lambda$-sausage body, as defined in Definition 1.3.
@@
Proof of Theorem 1.3: Substituting $\rho=\coth^{-1}(\lambda)$ in equation (3.6), we get that the area of the thick $\lambda$-sausage is
$|S|=\frac{|\partial S|}{\lambda}+2\pi\left(1-\sqrt{1-\frac{1}{\lambda^{2}}}\right).$ (3.7)
Therefore, for the area of a thick $\lambda$-concave body $K$ satisfies
$|K|\geq\frac{|\partial K|}{\lambda}+2\pi\left(1-\sqrt{1-\frac{1}{\lambda^{2}}}\right).$ (3.8)
Applying the appropriate scaling, when $K\subset\mathbb{H}^{2}(-c^{2})$, (3.8) becomes
@@
It is easy to check (applying the L’Hôpital Rule twice) that as $c\rightarrow 0$ from the left, equation (3.7) approaches the result in the Euclidean space proved in [4]. This concludes the proof of one direction of Theorem 1.3. The other direction is trivial.